Cleaning Price per Unit $ 360 640 1,500 Variable Cost per Unit $200 620 860 Units Sold per Year 9,000 900 100 Filling CappingA. is 801,600. I need B and C.
Answer #1

Painless Dental Clinic, Inc.

B) Break even Analysis:

Sales mix,

Cleaning = 9,000/10,000 = 90%

Filling = 900/10,000 = 9%

Capping = 100/10,000 = 1%

Weighted average contribution per unit,

= Contribution per unit × Sales mix

Selling price Variable cost Contribution per unit (C) Sales mix (S) (C × S)
Cleaning 360 200 160 90% 144
Filling 640 620 20 9% 1.8
Capping 1500 860 640 1% 6.4
152.2

Break even point = Fixed cost / Weighted average contribution per unit

= $ 520,000 / $ 152.2 per unit = 3,417 unit (approximately)

Break even sales revenue:

Break even units × Sales Mix × Selling price
Cleaning 3,417 units × 90% × $ 360 $ 11,07,108
Filling 3,417 units × 9% × $ 640 $ 196,819
Capping 3,417 units × 1% × $ 1,500 $ 51,255
$ 13,55,182

c) Sales revenue, earn $ 164,00 per year after taxes ($ 205,000 before taxes) (= 164,000/0.8)

Selling units = (Fixed cost + Target Income) / Weighted average contribution per unit

= (520,000+205,000)/152.2

= 725,000/152.2 = 4,763 units (approximately)

Sales revenue:

units × Sales mix × Selling price
Cleaning 4,763 units × 90% × $ 360 $ 15,43,212
Filling 4,763 units × 9% × $ 640 $ 2,74,349
Capping 4,763 units × 1% × $ 1500 $ 71,445
$ 18,89,006